# Depth to Anatomy: Learning Internal Organ Locations from Surface Depth Images

### v1

@@ Abstract
Automated patient positioning plays an important role in optimizing scanning procedure and improving patient throughput. Leveraging depth information captured by RGB-D cameras presents a promising approach for estimating internal organ positions, thereby enabling more accurate and efficient positioning. In this work, we propose a learning-based framework that directly predicts the 3D locations and shapes of multiple internal organs from single 2D depth images of the body surface. Utilizing a large-scale dataset of full-body MRI scans, we synthesize depth images paired with corresponding anatomical segmentations to train a unified convolutional neural network architecture. Our method accurately localizes a diverse set of anatomical structures, including bones and soft tissues, without requiring explicit surface reconstruction. Experimental results demonstrate the potential of integrating depth sensors into radiology workflows to streamline scanning procedures and enhance patient experience through automated patient positioning. The implementation and pretrained models are publicly available at https://github.com/EytanKats/orgloc.
@@
Radiology workflow, Organs localization
@@ 1 Introduction
The growing demand for radiology examinations underscores the need to optimize scanning procedures to reduce patient wait times and improve operational efficiency. A critical, yet time-intensive, aspect of this process is the patient preparation and planning phase, which is essential for optimizing image quality and diagnostic accuracy (van2020cinderellas). This phase typically involves manual positioning of the patient followed by scout imaging to facilitate detailed geometric planning for subsequent diagnostic scans. In standard clinical practice, a radiologic technologist initially aligns the patient on the table, often relying on external anatomical landmarks or laser positioning guides to approximate the region of interest. Once the patient is positioned inside the bore, a set of low-resolution, multi-planar scout images is acquired. These images are then used to visually confirm the patient’s position and to manually plan the geometry of the diagnostic sequences, such as slice orientation, field-of-view, and scan range (koken2009towards).
@@
Figure 1: Overview of the proposed automated patient positioning workflow. A depth sensor installed above the scanner bore captures a depth image of the patient lying on the table. This image is processed by a trained model that predicts the 3D bounding box of the target organ. The scanner table is then automatically adjusted to align the organ of interest with the center of the imaging field-of-view, reducing the need for manual repositioning and scout scans.
@@
Among the emerging technological approaches in this area, RGB-D camera-based systems have shown notable potential in enhancing patient setup and positioning (booij2021automated; booij2019accuracy). These integrated systems use depth-sensing to generate surface maps of the patient’s body. By analyzing these surface contours, it becomes possible to detect and localize key anatomical landmarks, which serve as proxies for internal anatomical regions of interest. This information can then be used to infer the optimal imaging volume and adjust the patient table position automatically prior to scan acquisition, without the need for scout imaging. For example, incetan2020rgb demonstrated the utility of RGB-D sensors in automatically localizing target regions such as the head and thorax with high spatial accuracy. Similarly, karanam2020towards employed depth-aware landmark detection models to support robust, contactless patient positioning in MRI and CT workflows.
In the proposed workflow, a depth sensor is integrated into the room above the scanner bore to capture a depth image of the patient lying on the table. This depth image is then processed by the trained model, which outputs the 3D coordinates of the bounding box enclosing the organ of interest. Based on this prediction, the scanner table can be automatically adjusted to ensure that the relevant anatomy is optimally centered within the imaging field-of-view, minimizing the need for manual repositioning or iterative scout imaging (Figure 1). This approach offers a contactless, fast, and standardized method of patient localization, which could be particularly beneficial in high-throughput clinical environments or for procedures requiring precise anatomical targeting.
To ensure robust performance across a wide range of anatomical variations and body types, we train our model on a large-scale dataset consisting of approximately 10,000 whole-body MRI scans from the German National Cohort (NAKO) study (bamberg2022whole). This dataset includes subjects of diverse ages, body sizes, and physiological characteristics, enabling the model to generalize effectively across different patient populations. Our experimental evaluation demonstrates the effectiveness of the proposed method in accurately localizing 41 distinct internal anatomical structures, including bones and soft tissue organs, using only depth images.
The key contributions of this work are as follows:
@@
    We demonstrate the feasibility and effectiveness of a learning-based framework for estimating 3D locations of internal anatomical structures solely from depth images of the patient’s body surface.
@@
    We introduce a convolutional neural network architecture, Pixel-to-Voxel (Pix2Vox), which maps 2D depth images to 3D volumetric segmentation masks, enabling organ localization from a single surface view.
@@
    We validate the generalizability of our model across a wide range of body types and anatomical variations by training and evaluating on a large-scale, population-level dataset comprising approximately 10,000 whole-body MRI scans.
2 Related Works
The prediction of internal anatomical structures from external body surface representations has gained increasing attention due to its potential to enable non-invasive, efficient, and patient-specific modeling. Early work by wu2018towards introduced a learning-based framework that generates volumetric anatomical phantoms of skeletal and pulmonary structures from triangulated meshes of the patient’s external surface. This study demonstrated the feasibility of inferring detailed 3D internal anatomy from external geometry using deep models trained on co-registered body surface and internal organ data.
Intermediate 3D representations for internal anatomy estimation.
Several approaches address the prediction of internal anatomy by first deriving an intermediate 3D representation from depth images, which is then used to infer internal structures. Building on the integration of statistical human body models with internal anatomy, recent research has extended the widely adopted Skinned Multi-Person Linear (SMPL) model (loper2015smpl) to incorporate internal organ representations. The original SMPL is a parametric, skinned 3D body model that encodes human pose and shape with a low-dimensional set of learned parameters, enabling the synthesis of realistic body meshes through linear blend skinning. Guo et al. (guo2022smpl) proposed SMPL-A, which augments SMPL with deformable internal organ geometries, allowing organ shapes to adapt dynamically with body pose. While effective in modeling pose-aware internal anatomy, SMPL-A requires a patient-specific full-body mesh to initialize internal organ shapes, limiting its generalizability and real-time applicability in clinical environments. Similarly, OSSO (keller2022osso) and BOSS (shetty2023boss) infer internal organ shapes and locations by optimizing anatomical plausibility within the SMPL framework based on external body surface meshes. Extending this paradigm, the HIT (Human Internal implicit Tissues) framework (keller2024hit) employs a deep implicit function to predict continuous volumetric tissue distributions conditioned on a parametric body surface mesh. HIT leverages large-scale synthetic datasets and a canonical space aligned with SMPL to infer anatomically coherent internal structures. Despite promising results, these SMPL-based approaches depend on accurate surface mesh reconstruction and template fitting, which are computationally intensive. The LOOC framework (henrich2025looc) introduces an implicit volumetric model to infer internal organ locations and approximate shapes from depth-derived surface point clouds.
Prediction of 2D organ projections.
Recent studies by geissler2025predicting and kats2025internal successfully utilize convolutional neural networks (CNNs) that operate directly on depth maps to predict 2D projections of organs in the coronal plane. These works demonstrate that CNNs can effectively extract spatial cues from external body surfaces to infer accurate organ locations.
Prediction of 3D features from 2D depth images.
Teixeira et al. (teixeira2023automated) proposed a CNN-based method that directly estimates critical scan planning parameters—such as lung centroids and patient isocenter—from raw depth images. This approach bypasses SMPL-based registration and operates solely on depth data, providing a lightweight and robust solution for specific imaging tasks such as isocenter calculation in thoracic MRI.
In contrast to previous work on organ localization, our method introduces a lightweight, end-to-end framework that directly predicts explicit 3D bounding boxes and estimates shapes for 41 internal anatomical structures from a single depth image. Leveraging a unified Pix2Vox convolutional architecture, our approach eliminates the need for intermediate steps such as 3D surface reconstruction or canonical space registration, enabling rapid inference and practical integration into clinical workflows. To foster reproducibility and facilitate further research, our code and pretrained models are publicly available at https://github.com/EytanKats/orgloc.
@@ 3 Materials and Methods
In this study, we leverage 10,020 full-body MRI scans from the National Cohort Study of Germany (NAKO) dataset (bamberg2022whole). This large and diverse dataset enables us to capture a broad spectrum of anatomical variability across age, sex, and body habitus, providing a robust foundation for training generalizable models of internal anatomy. Since paired depth images and organ segmentations are not natively available in the dataset, we synthetically generate training pairs from the MRI data.
To create depth images, we simulate depth sensor output by rendering the outer contours of the patient’s body from MRI volumes, mimicking the top-down perspective of real-world RGB-D sensors typically mounted above the scanner table. Simultaneously, we apply state-of-the-art automated segmentation tools to extract ground-truth 3D labels for 41 internal anatomical structures, including bones and soft tissue organs (Section 3.1).
The resulting dataset, composed of 2D depth images paired with corresponding 3D anatomical segmentations, serves as the training input for our proposed learning pipeline. We devised CNN and train Pix2Vox convolutional neural network, which maps each 2D depth image to a 3D volumetric prediction of the target organ bounding boxes Section 3.2).
3.1 Data Preparation
To simulate depth images from MRI volumes, we begin by normalizing each 3D MRI scan to the intensity range $[0,1]$. Next, we apply a global intensity threshold of 0.02 to generate a binary mask that captures the patient’s body. Small noise components and isolated artifacts are subsequently removed using binary morphological opening with a spherical structuring element, producing a cleaner mask. We then extract 2D orthographic depth projections in the coronal plane, simulating the top-down view of a depth sensor positioned above the patient. Specifically, for each pixel in the image plane, the depth value is computed as the closest point of intersection along the anterior-posterior axis with the segmented patient’s body. The resulting depth maps are normalized again to the range $[0,1]$. To suppress non-anatomical structures such as the patient table and background noise, we apply a depth cutoff, setting all pixel values below 0.3 to 0. This effectively removes table artifacts and regions far below the simulated camera plane. Finally, to refine the depth image, we apply grayscale morphological opening, which smooths remaining artifacts while preserving significant anatomical features such as limb contours and body curvature. The resulting depth images offer a standardized approximation of the patient’s body surface from a fixed sensor viewpoint and serve as input to our learning framework.
To generate high-quality anatomical reference labels, we leveraged recent state-of-the-art automated whole-body MRI segmentation tools, including TotalSegmentatorMRI (d2024totalsegmentator), MRSegmentator (hantze2024mrsegmentator), and TotalVibeSegmentator (graf2024totalvibesegmentator). These models, trained on large-scale MRI datasets, are capable of accurately segmenting a wide range of anatomical structures with high spatial precision. To further increase the number of segmented labels, the MR volumes were converted into synthetic CT volumes to utilize segmentation models available for CT data (wasserthal2023totalsegmentator; jaus2023towards), following the approach described in mensing2025mr. An aggregation procedure was implemented to select, for each anatomical structure, the segmentation mask with the highest accuracy. When a structure was available from multiple models, MR-based segmentations were prioritized over CT-based ones. The MRI segmentation models were ranked according to the similarity of their training data to our dataset, in the following order: TotalVibeSegmentator, MRSegmentator, and TotalSegmentatorMRI. For structures exclusively available from CT-based models, TotalSegmentator was preferred over Atlas DAP. The resulting aggregated masks encompass 137 distinct structures, including organs, vessels, bones, and tissue types.
Given the varying quality of automated segmentations across different structures and the need for relevance to clinical workflow optimization, we performed a careful selection process. In consultation with domain experts, and through extensive visual inspection of the output masks, we selected 41 anatomical structures (Table 1). These included bones and soft-tissue organs that are both (1) of practical importance for guiding scan planning and patient positioning, and (2) reliably segmented across a broad range of patients. To further improve the consistency of the labels and reduce artifacts introduced by the segmentation models, we applied a series of post-processing steps. These included hole filling to restore anatomical completeness, and connected components analysis to eliminate spurious isolated regions or misclassified voxels. These refinements ensured cleaner and more robust ground-truth volumes for training the model.
Figure 3: Mean absolute Detection Offset Error (DOE) averaged across all organs for each bounding box side (left-right, anterior-posterior, superior-inferior). The CNN models demonstrates strong localization accuracy, particularly in the coronal (left-right) and vertical (superior-inferior) dimensions, significantly outperforming the mean model baseline. They achieve sub-10mm average error in the coronal plane. Performance degrades along the anterior-posterior axis due to limited depth cues available from the coronal input image. Standard deviation is depicted as error bars extending in the positive direction only.

### v2

@@ Abstract
Automated patient positioning can improve radiology workflow efficiency by reducing the time required for manual table adjustments and scout-based scan planning. We propose a learning-based framework that predicts 3D organ locations and shapes for 41 anatomical structures, including both bones and soft tissues, directly from a single 2D depth image of the body surface. Leveraging $10,020$ whole-body MRI scans from the German National Cohort (NAKO) dataset, we synthetically generate depth images paired with anatomical segmentations to train a convolutional neural network for volumetric organ prediction. Our method achieves a mean dice similarity coefficient of $0.44\pm 0.2$ and and a symmetric average surface distance of $7.69\pm 5.68$ mm across all structures. Furthermore, the model derives organ bounding boxes with a mean absolute detection offset of $10.99\pm 5.54$ mm. Qualitative results on real-world depth images confirm the ability of the model to generalize to practical clinical settings. These findings suggest that depth-only organ localization can support automated patient positioning reducing setup time, minimizing operator variability, and improving patient comfort. The implementation and pretrained models are publicly available at https://github.com/EytanKats/orgloc.
@@
Radiology workflow automation, Automated patient table positioning, Body surface-to-3D anatomy prediction, Hybrid 2D–to-3D neural architecture.
@@ 1 Introduction
The growing demand for radiological examinations highlights the need to optimize scanning workflows in order to reduce patient waiting times and improve operational efficiency. An important yet time-consuming component of this workflow is the scan preparation and planning phase, which plays a crucial role in ensuring image quality and diagnostic reliability (van2020cinderellas). According to (streit2021analysis), scan preparation requires on average 5.7 minutes, corresponding to 11% of the total examination time. This phase typically involves manual patient positioning followed by scout imaging to enable geometric planning of the subsequent diagnostic acquisition. In routine clinical practice, a radiologic technologist first aligns the patient on the scanner table, often relying on external anatomical landmarks and laser positioning guides to approximate the region of interest. After positioning inside the scanner bore, low-resolution multi-planar scout images are acquired. These images are then used to verify patient positioning and to manually define the geometry of the diagnostic sequences, including slice orientation, field of view, and scan range (koken2009towards).
@@
1.1 Related Works
Surface-based patient positioning systems.
@@ 1.1 Related Works
Among the emerging technological approaches for automating patient table positioning, RGB-D camera-based systems have demonstrated significant potential for enhancing clinical workflows (booij2021automated; booij2019accuracy). These systems utilize depth sensing to capture the external body surface, identifying anatomical landmarks that serve as proxies for internal regions of interest. This information enables the localization of key structures, which can be used to infer the optimal imaging volume and adjust the patient table automatically prior to scan acquisition, potentially eliminating the need for scout imaging. For instance, incetan2020rgb demonstrated the utility of RGB-D sensors in localizing regions such as the head and thorax with high spatial accuracy. Similarly, karanam2020towards employed depth-aware landmark detection to support robust, contactless positioning in MRI and CT workflows. However, these approaches primarily focus on external positioning and do not explicitly infer the 3D locations or shapes of internal anatomy.
Parametric body-model-based internal anatomy estimation.
Several approaches address the prediction of internal anatomy by first deriving an intermediate 3D representation of the body, which is then used to infer internal structures. Building on the integration of statistical human body models with internal anatomy, recent research has extended the widely adopted Skinned Multi-Person Linear (SMPL) model (loper2015smpl) to incorporate internal organ representations. The original SMPL is a parametric, skinned 3D body model that encodes human pose and shape with a low-dimensional set of learned parameters, enabling the synthesis of realistic body meshes through linear blend skinning. Guo et al. (guo2022smpl) proposed SMPL-A, which augments SMPL with deformable internal organ geometries, allowing organ shapes to adapt dynamically with body pose. While effective in modeling pose-aware internal anatomy, SMPL-A requires a patient-specific full-body mesh to initialize internal organ shapes, limiting its generalizability and applicability in clinical environments. Similarly, OSSO (keller2022osso) and BOSS (shetty2023boss) infer internal organ shapes and locations by optimizing anatomical plausibility within the SMPL framework based on external body surface meshes. However, these methods are constrained by a limited scope: OSSO focuses exclusively on the skeletal system, and BOSS models only 12 specific anatomical structures. Extending this paradigm, the HIT (Human Internal implicit Tissues) framework (keller2024hit) employs a deep implicit function to predict continuous volumetric tissue distributions conditioned on a parametric body surface mesh. While HIT achieves anatomically coherent internal structures, it focuses on general tissue distributions rather than the localization of discrete, individual organs. Furthermore, SMPL-based approaches typically require a complete 3D surface mesh of the entire body for initialization or template registration which are not available in clinical settings.
Estimation of internal anatomy from depth images.
Recent studies by geissler2025predicting and kats2025internal demonstrate that convolutional neural networks operating directly on depth maps can infer three-dimensional organ locations and approximate shapes from external body surface information. These works highlight that depth-derived surface cues contain sufficient anatomical information to estimate internal structures without explicit geometric reconstruction. While these works demonstrate that external body surface information contains useful cues about internal anatomy, their predictions are limited to planar projections rather than full volumetric organ localization. Similarly, teixeira2023automated proposed a CNN-based method that estimates specific scan-planning parameters, such as lung centroids and patient isocenter, directly from depth images. Although effective for targeted task, such predictions provide only sparse anatomical descriptors and do not capture full organ extent or shape.
A closely related concurrent study is the LOOC framework (henrich2025looc), which infers internal organ locations and approximate shapes using an implicit volumetric occupancy model conditioned on depth-derived surface point clouds. The approach first reconstructs a patient-specific external surface representation from depth data and subsequently predicts internal anatomy through a learned continuous occupancy field, enabling anatomically plausible volumetric estimates.
1.2 Contribution
This work presents a learning-based framework for estimating internal volumetric anatomical structures directly from a single anterior depth image. In contrast to surface-based patient positioning systems, which rely on external landmark detection or approximate region-of-interest estimation from body surface data (booij2021automated; booij2019accuracy; incetan2020rgb; karanam2020towards), our method explicitly predicts the three-dimensional locations and approximate shapes of internal anatomical structures. Unlike parametric body-model approaches that infer internal anatomy through template fitting or full-body surface mesh reconstruction within statistical body models (guo2022smpl; keller2022osso; shetty2023boss; keller2024hit), the proposed framework performs direct end-to-end depth-to-volume inference without intermediate body model fitting. Furthermore, whereas several recent depth-based approaches focus on predicting sparse anatomical descriptors such as organ centroids, isocenters, or limited projection-based organ localization (teixeira2023automated; geissler2025predicting; kats2025internal), our method produces explicit volumetric organ segmentations suitable for anatomically informed scan planning. Compared to the closely related concurrent LOOC framework (henrich2025looc), which employs an implicit volumetric occupancy model conditioned on reconstructed surface point clouds, our approach differs substantially in methodology by directly predicting volumetric anatomy from depth images using a hybrid 2D-to-3D convolutional neural network. To support reproducibility and further research, the code and pretrained models are publicly available at https://github.com/EytanKats/orgloc.
The key contributions of this work are summarized as follows:
@@ 1.2 Contribution
    We introduce a hybrid 2D-to-3D convolutional neural network architecture (Pix2Vox) that maps a single anterior depth image to volumetric segmentations of 41 anatomical structures, providing explicit spatial information for automated radiological positioning.
@@
    We demonstrate large-scale training and validation on a dataset comprising 10,020 MRI-derived paired depth and volumetric anatomy samples, enabling robust modeling of anatomical variability across a diverse population.
@@
    We qualitatively show that a model trained purely on synthetic depth–anatomy pairs generalizes to real-world depth images of volunteers positioned in an MRI scanner, including scenarios with clothing, supporting the feasibility of practical clinical deployment.
@@ 2 Materials and Methods
In this study, we leverage $10,020$ whole-body MRI scans from the German National Cohort (NAKO) dataset (bamberg2022whole). This diverse dataset captures a broad spectrum of anatomical variability across age, sex, and body habitus, providing a foundation for training generalizable models of internal anatomy. As paired depth images and expert organ segmentations are not natively provided, we designed a pipeline to synthetically generate supervised training pairs directly from the MRI volumes.
To synthesize the input data, we simulate depth sensor acquisitions by rendering the anterior contours of the patient’s body from the MRI volumes. This process mimics the orthographic top-down perspective of a sensor mounted above the scanner table (Section 2.2). To extract ground-truth 3D masks for 41 internal anatomical structures we utilize the TotalSegmentatorMRI tool (d2024totalsegmentator)(Section 2.3).
The resulting dataset, comprising 2D depth images paired with 3D anatomical segmentations, is used to train a Pix2Vox convolutional neural network (Section 2.5). This architecture is optimized to map 2D surface depth information to 3D volumetric predictions of the underlying target organs. Once trained, the model enables a depth sensor-augmented radiology workflow, facilitating automated patient positioning through anatomical structures localization (Section 2.1).
2.1 Radiology Workflow with Integrated Depth Sensor
Figure 1: Overview of the proposed automated patient positioning workflow. A depth sensor mounted above the bore captures the patient’s surface geometry. Through a spatial calibration and coordinate transformation process, the depth data is mapped into the scanner’s coordinate system. The trained model predicts 3D organ segmentation within the scanner coordinate system. From this, a bounding box is derived to enable automated table translation, aligning the target anatomy with the imaging isocenter.
In the proposed workflow, a depth sensor is mounted above the scanner bore to capture an anterior depth image of the patient on the table. The sensor is spatially calibrated with the MRI system, allowing the acquired depth data to be transformed into a surface point cloud within the scanner’s coordinate frame (booij2021automated; booij2019accuracy). This point cloud is subsequently rendered as an orthographic depth projection (Section 2.4). The resulting processed image is then fed into the trained model, which predicts the 3D segmentations of the anatomical structures of interest in scanner space. Based on these predictions, the scanner table can be automatically translated to ensure the target anatomy is optimally centered within the imaging isocenter, thereby minimizing the reliance on manual adjustments or iterative scout imaging (Figure 1). This approach offers a contactless, fast, and standardized method of patient localization, which could be particularly beneficial in high-throughput clinical environments.
@@ 2.1 Radiology Workflow with Integrated Depth Sensor
2.2 Synthetic Depth Image Generation
To facilitate the training of a body surface-to-internal organs model, we designed a deterministic pipeline to synthesize 2D orthographic depth projections from 3D MRI volumes. The process transforms volumetric data $\mathcal{V}\in\mathbb{R}^{H\times W\times D}$ into depth maps $D(x,z)$
First, MRI intensity values are normalized to the range $[0,1]$. A binary mask $M$ is generated by applying a low-intensity threshold $\epsilon=0.02$ to reduce the inherent noise:
$M_{i,j,k}=[\mathcal{V}_{i,j,k}>0.02]$ (1)
To remove floating artifacts and ensure structural continuity of the skin surface, we apply binary morphological opening $\gamma_{B}(M)$ using a spherical structuring element with diameter equal to 3.
The depth image $D$ is generated via a ray-casting approach along the anterior-posterior axis. For each lateral and longitudinal coordinate $(x,z)$, the distance to the first encountered foreground voxel is computed:
$d(x,z)=\text{argmax}_{y}(M_{x,y,z})$ (2)
The raw indices are normalized by the maximum volume height $H_{max}$ and inverted so that intensity values are inversely proportional to the distance from the sensor:
$D_{norm}(x,z)=1-\frac{d(x,z)}{H_{max}}\quad\text{for }d(x,z)>0$ (3)
To clean signal from the scanner table and hardware artifacts, we apply a clipping threshold, zeroing all pixels with $D_{norm}<0.3$. High-frequency surface noise and acquisition spikes are suppressed using grayscale morphological opening with an $11\times 11$ kernel.
To bridge the domain gap between clean synthetic depth maps and real-world clinical acquisitions, we implemented a stochastic augmentation pipeline designed to simulate the depth profile of hospital gowns, blankets, and clothing. The augmented depth map $D_{aug}$ is generated by modifying the synthetic depth $D$ within the patient silhouette mask $M$ as follows:
$D_{aug}=G_{\sigma}\left(D+[B_{lines}+B_{clusters}]\cdot M\right)$ (4)
where $G_{\sigma}$ denotes a global Gaussian smoothing kernel with $\sigma=5$, $B_{lines}$ is a line-like fold simulation and $B_{clusters}$ cluster-like noise simulation.
Line-like Fold Simulation ($B_{lines}$). We simulate fabric creases by generating $N\in[5,20]$ random linear segments. For each segment, the orientation is sampled from $\mathcal{U}(0,2\pi)$ and the length from $\mathcal{U}(H/12,H/2)$, where $H$ is the image height. Each segment is assigned a randomized thickness of 141-4 pixels and blurred using a Gaussian kernel with $\sigma_{line}\in[S/20,S/5]$, where $S=100$ is a spatial scale factor. Folds are assigned a random polarity $p\in\{-1,1\}$ and scaled to a maximum displacement of $30$ mm.
Cluster-like Variations ($B_{clusters}$). To mimic the effect of bunched blankets or hospital gowns, we generate multi-scale, low-frequency noise clusters. This is achieved by sampling random noise at three spatial scales $\{0.5S,S,2.0S\}$. For each scale, a coarse noise grid is generated and then expanded to the full image resolution using cubic interpolation. Each upscaled layer is refined with a Gaussian smoothing kernel where the standard deviation is scaled to the specific layer’s frequency ($\sigma=s/10$). This process produces a continuous, bias field that simulates bulky surface irregularities. The integrated cluster bias is scaled to a maximum displacement of $30$ mm.
The augmentation process ensures that the synthetic training data exhibits the non-rigid, low-frequency spatial variations characteristic of clothes and blankets, forcing the network to prioritize global anatomical proportions over local surface textures.
2.3 Generation of Volumetric Ground Truth Labels
To generate anatomical reference labels, we utilized TotalSegmentatorMRI (d2024totalsegmentator), a robust deep learning framework capable of segmenting 56 distinct anatomical structures in whole-body MRI. To ensure the reliability of these automated segmentations as ground-truth data, we implemented a post-processing pipeline designed to improve topological consistency and eliminate segmentation artifacts. Specifically, we applied morphological hole filling to ensure anatomical completeness and 3D connected component analysis (CCA) to remove small isolated, spurious voxel clusters and misclassified regions.
Due to the significant memory requirements of the 3D convolutional decoder we refined the label set to 41 target structures (Table 1). This selection was based on a dual-criteria approach: (1) a qualitative visual assessment of the automated segmentation accuracy across the cohort, and (2) the clinical relevance of the structure for automated scan planning and patient positioning. Structures with lower segmentation reliability or minimal impact on clinical workflow were excluded to optimize the model’s computational efficiency.
Figure 3: Mean absolute Detection Offset Error (DOE) averaged across all organs for each bounding box side (left-right, anterior-posterior, superior-inferior).
2.4 Processing of Real-World Depth Images
To evaluate the robustness of our proposed model, we qualitatively assess its performance on real-world depth maps acquired from human volunteers using a Microsoft Kinect v2 sensor (Section 3.4.3). Raw depth frames $\mathcal{Z}\in\mathbb{R}^{U\times V}$ are first processed using a spatial Region-of-Interest (ROI) filter to isolate the subject from the scanner environment. This filter retains voxels within a valid depth range of $[1000,2000]$ and a lateral bounding box defined by $u\in[450,1050]$ and $v\in[400,700]$. Utilizing the intrinsic parameters of the depth sensor and extrinsic calibration parameters, the depth coordinates are transformed into the coordinate system of the scanner, yielding a sparse point cloud $\mathbf{P}=\{x,y,z\}$.
The final anterior orthographic depth image is generated by raycasting $\mathbf{P}$ onto a 2D grid. To account for no-data pixels we employ linear interpolation to reconstruct a continuous surface. The resulting depth values are min-max normalized and contrast-inverted:
$D_{real}=1.0-\frac{Z-Z_{min}}{Z_{max}-Z_{min}}$ (5)
This transformation ensures that the region in closest proximity to the sensor is represented by high-intensity values, maintaining distribution consistency with the synthetic training data $D_{norm}$ described in Section 2.2.